# ISP-AD: A Large-Scale Real-World Dataset for Advancing Industrial Anomaly Detection with Synthetic and Real Defects

### v2

@@ Abstract
Automatic visual inspection using machine learning-based methods plays a key role in achieving zero-defect policies in industry. Research on anomaly detection approaches is constrained by the availability of datasets that represent complex defect appearances and imperfect imaging conditions, which are typical to industrial processes. Recent benchmarks indicate that most publicly available datasets are biased towards optimal imaging conditions, leading to an overestimation of the methods’ applicability to real-world industrial scenarios. To address this gap, we introduce the Industrial Screen Printing Anomaly Detection dataset (ISP-AD). It presents challenging small and weakly contrasted surface defects embedded within structured patterns exhibiting high permitted design variability. To the best of our knowledge, it is the largest publicly available industrial dataset to date, including both synthetic and real defects collected directly from the factory floor. In addition to the evaluation of defect detection performance of recent unsupervised anomaly detection methods, experiments on a mixed supervised training approach, incorporating both synthesized and real defects, were conducted. Even small amounts of injected real defects prove beneficial for model generalization. Furthermore, starting from training on purely synthetic defects, emerging real defective samples can be efficiently integrated into subsequent scalable training. Research findings indicate that supervision by means of both synthetic and accumulated real defects can complement each other, meeting demanded industrial inspection requirements such as low false positive rates and high recall. The presented unsupervised and supervised dataset splits are designed to emphasize research on unsupervised, self-supervised, and supervised approaches, enhancing their applicability to industrial settings.
@@ 1.1 Contributions
The aim of this publication is twofold. First, to bridge the gap to industry, we introduce a dataset derived from a real-world industrial use case in screen printing. This dataset provides both unsupervised and supervised training data, comprising synthetic and real defects collected during production. Second, we investigate a mixed supervised training approach that efficiently utilizes available weakly labeled data consisting of both, synthetic and real defects. By means of this publicly available data set, further research in leveraging synthetic as well as real defective data for industrial anomaly detection should be emphasized.
Thus, the main contributions can be summarized to:
@@
  • Introduction of a novel dataset of structured patterns, captured using three different optical modalities. The dataset originates from a real-world industrial use case in screen printing, including permitted process-specific data variability. With a total of 312 674312674312\,674312 674 fault-free samples and 246 375246375246\,375246 375 defective samples, where 245 664245664245\,664245 664 are synthetic and 711 are real, it represents the largest industrial defect detection dataset to date, enabling both unsupervised and supervised training scenarios.
@@
  • Image and pixel-level benchmarks of SOTA unsupervised approaches on the proposed demanding imaging modalities, containing small and weakly contrasted defects with high permitted design variability.
  • The mixed supervised training procedure introduced by [26] was further investigated by utilizing various proportions of injected real defects. Starting from purely synthetic training, the incremental incorporation of accumulated weakly labeled real defects led to a notable improvement in defect detection performance, particularly for challenging small and weakly contrasted defects. Additionally, the limited use of previously unseen defects significantly enhanced generalization performance, complementing synthetic training data. Thus, emerging defective samples with previously unseen feature distributions can be efficiently integrated into the mixed training procedure.
  • Novel industrial benchmark, enabling research on unsupervised, self-supervised and supervised anomaly detection approaches.
@@ 2.1 Supervision in Industrial Anomaly Detection
In contrast to sampling from the examined data distribution, DRAEM [74] generates diverse defect shapes by extracting textures from different domains (out-of-distribution) [17] using masks generated via Perlin noise [50]. Additional approaches use gaussian or simplex noise [51] added onto normal images to generate synthetic defects [67, 13]. More realistic defects, referred to as in-distribution defects, can be achieved by generating defect textures on normal samples by means of random walks [25].
@@ 2.2 Industrial Anomaly Detection Datasets
Table 1: Comparison of SOTA anomaly detection datasets by sample size, defect data availability, and generation method (manual, synthetic, real-world). The proposed dataset, ISP-AD, introduces supervised and unsupervised data splits with synthetic and real defects, collected at the factory floor.
@@
VisA 10 8211082110\,82110 821 $9621$ $1200$ 0 12 rgb manually
Real-IAD 151 050151050151\,050151 050 99 7219972199\,72199 721 51 3295132951\,32951 329 0 30 rgb manually
@@
DAGM 16 1001610016\,10016 100 14 0001400014\,00014 000 2100 1050 10 mono synthetic
MIAD 105 000105000105\,000105 000 87 5008750087\,50087 500 17 5001750017\,50017 500 0 7 rgb synthetic
@@
VAD $5000$ $3000$ 2000 $0001$ rgb real-world
ISP-AD 559 049559049559\,049559 049 312 674312674312\,674312 674 711/245 664245664245\,664245 664 446 3 rgb/mono real-world/synthetic
@@
Both DAGM [46] and MIAD [5] are based on synthetic defect generation. DAGM, published in 2007, contains 10 synthetically generated classes summing up to 14 0001400014\,00014 000 fault-free background textures and 2100 defectives, one for each background texture. In contrast, the more recent MIAD contains 105 000105000105\,000105 000 images of various outdoor industrial maintenance scenarios. MIAD leverages 3D graphics software [19] to render realistic 3D scenes with varying surface textures, backgrounds and viewpoints on both fault-free and defective objects.
@@ 3 Industrial Screen Printing Anomaly Detection Dataset: ISP-AD
    High permitted sample variance arising from the sample itself, preprocessing, and imperfect imaging conditions.
@@ 3.1 Industrial Screen Printing and Defect Classes
[…]
@@ 3.2 Dataset Generation
The generation of appropriate datasets builds the basis for achieving the required defect detection performance of machine learning based methods applied on the production line. Therefore, the design of validation and test splits mimicking the inspection process with its imbalanced data distributions is of utmost importance to perform reliable evaluations. Furthermore, generated data distributions must resemble the imperfect imaging conditions caused by environmental influences (e.g., varying illumination), data processing (varying background, viewpoint and segmentation artifacts), as well as permitted sample variance.
The proposed dataset is based on the work presented in [35], including its efficient data preprocessing and labeling workflow. Prior to image acquisition, fault-free and defective samples were pre-sorted by domain experts. The sample state is characterized as fixed, thus additional temporal sample alterations can be excluded. Moreover, the samples were selected from different production batches, resembling the permitted design variability. Images were acquired by means of all three optical modalities (LSM-1, LSM-2, ASM) of the inspection system demonstrator installed on the factory floor. The acquired images of fault-free and defective samples were assigned to separate training and test splits. This step avoids any unwanted correlation of fault-free and defective patches during the subsequent patch extraction process.
The extraction of augmented fault-free training patches (256 x 256 px) at random positions within the region of interest (ROI) is integrated into an automated procedure (steps 1 to 5 in section 5.1 [35]). Furthermore, following the central cropping in step 4, synthetic defects according to section 5.2 are generated in 50 % of the fault-free patches. The utilized algorithm [25] is based on a stochastic process resembling a random walk with momentum. Thus, the randomly generated defect textures enabled the synthetization of a wide range of punctual defect appearances, in all three optical modalities (Fig. 3). As a result, balanced supervised training datasets were obtained.
@@
Above described workflow minimizes the manual labeling effort to known defective samples and regions. The underlying image-level labels can be considered ’weak’ since no ground truth masks are provided, thereby eliminating the need for elaborate pixel-level annotation. In general, patch-based processing is an efficient augmentation technique in upscaling available data, enabling extraction of several 100 000100000100\,000100 000 patches from a few acquired samples with large FOVs.
@@ 3.2.1 Data Splits
The aim of this publication is to advance research in solving real-world industrial inspection problems by leveraging both synthetic and available real defects accumulated during production processes. Therefore, supervised and unsupervised datasets for each optical modality, acquired from the factory floor, have been created. The supervised training data splits include up to approximately 100 000100000100\,000100 000 augmented fault-free patches and 100 000100000100\,000100 000 synthetic defective patches. Additionally, real defective patches, extracted from defective samples, are categorized into two distinct defect groups: area and points. For unsupervised approaches, the fault-free training data includes up to 3678 patches in LSM-1, with reduced splits containing a maximum of 500 patches available for all modalities. These smaller splits are designed to minimize computational effort for e.g. memory demanding methods. The test data for both supervised and unsupervised approaches is identical, mimicking typical imbalanced inspection data distributions. Table 7 summarizes the proposed dataset with its synthetic and real defects available for supervision.
@@ 3.2.4 Additional Data
In addition to proposed data splits, supplementary fault-free patches for all modalities are provided. These patches are the unaugmented counterparts to the fault-free supervised training data described above. They consist of 20 0002000020\,00020 000 and 10 0001000010\,00010 000 fault-free patches in rgb format for LSM-1 and LSM-2, and 20 0002000020\,00020 000 patches in single-channel format for ASM, all stored as .png files.
@@ 4 Defect Detecion Methods
The recently published work [35] introduced an efficient data preprocessing workflow utilizing weakly labeled defective data in a supervised training approach. Furthermore, the possible extension to a mixed training approach using both synthetic and available real defective data [26] was outlined, enhancing defect detection sensitivity of overlooked defect classes.
Therefore, the evaluation of the mixed training approach on the presented demanding imaging modalities should give further insights in leveraging supervised methods in industrial anomaly detection. For comparison several SOTA unsupervised anomaly detection methods are investigated. Utilized approaches as well as their training settings will be described in the following section.
@@ 4.1 Mixed Supervised Training
The proposed mixed supervised training approach showed increased detection performance by leveraging both synthetic and real defects during training, compared to training separately on synthetic or real ones [26]. Thus, depending on the availability of real defects, the method can be applied in cold-start scenarios, starting with purely synthetic training and transitioning to a mixed setup as real defects accumulate. This is an important capability in industrial applications, as the model can be improved step by step with newly available defective samples.
@@
However, the ability to synthesize certain defect classes using the algorithm in [25], such as squeegee strokes shown in Fig. 2, is limited or even infeasible. Thus, even small amounts of previously unseen real defects may contribute to learning feature representations that could not be captured by synthetization or were absent in the available real data. Additionally, incorporating real defects, similar to synthesized ones, increases the feature diversity of already seen defects.
@@
Utilizing the above introduced datasets a DCNN, pretrained on ImageNet [32], is trained in a supervised manner. Synthetic defective patches were generated in 50 % of the fault-free patches. During training, each patch in this balanced stream is replaced by an augmented real defective patch with a probability of considerably less than 50 %. Thus, the objective is to complement the synthetic defects, not to replace them.
Supervised binary classification is inherently a closed-set problem, in contrast to the open-set problem of anomaly detection described above. Adapting the classifier to emerging, previously unseen defect classes with a minimal set of recently available samples would be an important capability in industrial settings.
@@
The following experiments in section 5.2 aim to investigate the generalization capabilies of the proposed mixed training approach for previously unseen defects (e.g. that were not able to be synthesized). In addition, its scalability is examined under varying proportions of collected real defects within all imaging modalities: LSM-1, LSM-2 and ASM.

### v3

@@ Abstract
Automatic visual inspection using machine learning plays a key role in achieving zero-defect policies in industry. Research on anomaly detection is constrained by the availability of datasets that capture complex defect appearances and imperfect imaging conditions, which are typical of production processes. Recent benchmarks indicate that most publicly available datasets are biased towards optimal imaging conditions, leading to an overestimation of their applicability in real-world industrial scenarios. To address this gap, we introduce the Industrial Screen Printing Anomaly Detection Dataset (ISP-AD). It presents challenging small and weakly contrasted surface defects embedded within structured patterns exhibiting high permitted design variability. To the best of our knowledge, it is the largest publicly available industrial dataset to date, including both synthetic and real defects collected directly from the factory floor. Beyond benchmarking recent unsupervised anomaly detection methods, experiments on a mixed supervised training strategy, incorporating both synthesized and real defects, were conducted. Experiments show that even a small amount of injected, weakly labeled real defects improves generalization. Furthermore, starting from training on purely synthetic defects, emerging real defective samples can be efficiently integrated into subsequent scalable training. Overall, our findings indicate that model-free synthetic defects can provide a cold-start baseline, whereas a small number of injected real defects refine the decision boundary for previously unseen defect characteristics, thereby meeting key industrial inspection requirements such as low false positive rates and high recall. The presented unsupervised and supervised dataset splits are designed to emphasize research on unsupervised, self-supervised, and supervised approaches, enhancing their applicability to industrial settings.
@@ 1.1 Contributions
The aim of this publication is twofold. First, to bridge the gap to industry, we introduce the Industrial Screen Printing Anomaly Detection Dataset (ISP-AD), derived from a real-world industrial use case in screen printing. This dataset provides both unsupervised and supervised training data, comprising synthetic and real defects collected during production. Second, we investigate a mixed supervised training strategy that efficiently utilizes available weakly labeled data consisting of both synthetic and real defects.
We envision the ISP-AD dataset as a research and development resource for advancing anomaly detection methods under realistic industrial conditions. Furthermore, we hope that it will contribute to a more application-oriented perspective, by helping to define paradigms and approaches that align with real-world requirements.
Thus, the main contributions of this work can be summarized as follows:
@@
  • Introduction of the Industrial Screen Printing Anomaly Detection Dataset (ISP-AD) : a novel large-scale dataset of structured patterns captured using three different optical modalities. The dataset originates from a real-world industrial screen printing process and includes permitted process-specific variability. With a total of 312 674312\,674 fault-free samples and 246 375246\,375 defective samples (of which 245 664245\,664 are synthetic and 711 are real), it represents the largest publicly available industrial defect detection dataset to date, enabling both unsupervised and supervised training scenarios (Future Applications).
@@
  • Benchmarking of SOTA unsupervised methods: comprehensive image- and pixel-level evaluations on ISP-AD highlight the challenges posed by small and weakly contrasted defects embedded within high permitted design variability, establishing ISP-AD as a demanding benchmark under industrial conditions.
  • Extended experimental investigation and formalization of mixed supervised training: building on our earlier work on mixed training [28], we analyze the incremental incorporation of weakly labeled real defects into synthetic cold-start training across multiple modalities and defect fractions. We formalize the strategy as a stochastic batch-level injection scheme, where real defects replace patches in a balanced synthetic stream with an injection probability $p_{\text{inj}}\in[1/B,0.25]$. This condition ensures that real defects complement rather than dominate the synthetic stream, providing controllable stochasticity. Experiments show that even small amounts of previously unseen defects can substantially enhance generalization, efficiently complementing synthetic training data.
  • Methodological insights into adaptive and scalable training strategies: the mixed supervised training is proposed as an adaptive closed-set strategy operating under a restricted open-set paradigm, where synthetic defects define an initial decision boundary and real defects are incorporated in subsequent retraining stages for iterative refinement within the known process domain. Ablations across ResNet [29], EfficientNet [72], and ConvNeXt [47] backbones confirm the architecture-agnostic scalability of the approach, demonstrating robustness under different accuracy–efficiency trade-offs and suitability for industrial deployment.
The structure of the research work is as follows: Section 2 gives an overview of emerging industrial anomaly detection methods as well as available datasets, and highlights existing research gaps. Section 3 introduces the real-world ISP-AD dataset, including the dataset generation process, available data splits, and its limitations. Section 4 describes the applied defect detection methods, the proposed mixed supervised training, and the investigated SOTA unsupervised methods. Section 5 presents the benchmark results across all three optical modalities. Finally, section 6 concludes the work and outlines future research directions.
@@ 2.1 Supervision in Industrial Anomaly Detection
In contrast to sampling from the examined data distribution, DRAEM [84] generates diverse defect shapes by extracting textures from different domains (out-of-distribution) [18] using masks generated via Perlin noise [55]. Additional approaches use gaussian or simplex noise [56] added onto normal images to generate synthetic anomalies [75, 14]. More realistic defects, referred to as in-distribution defects, can be achieved by generating defect textures on normal samples by means of random walks [27].
@@ 2.2 Industrial Anomaly Detection Datasets
Table 1: Comparison of SOTA anomaly detection datasets by sample size, defect data availability, and generation method (manual, synthetic, real-world). The proposed dataset, ISP-AD, introduces supervised and unsupervised data splits with synthetic and real defects, collected at the factory floor. In addition, ISP-AD addresses key challenges of industrial inspection that are underrepresented in existing datasets: (i) real-world imperfect imaging conditions and high permitted sample variations; (ii) imbalanced test data splits that reflect realistic inspection scenarios; and (iii) small and weakly contrasted defects on structured patterns relative to large sample areas.
@@
VisA 10 82110\,821 $9621$ $1200$ 0 12 rgb manually
Real-IAD 151 050151\,050 99 72199\,721 51 32951\,329 0 30 rgb manually
@@
DAGM 16 10016\,100 14 00014\,000 2100 1050 10 mono synthetic
MIAD 105 000105\,000 87 50087\,500 17 50017\,500 0 7 rgb synthetic
@@
VAD $5000$ $3000$ 2000 $1000$ 1 rgb real-world
ISP-AD 559 049559\,049 312 674312\,674 711/245 664245\,664 446 3 rgb/mono real-world/synthetic
@@
Both DAGM [51] and MIAD [5] are based on synthetic defect generation. DAGM, published in 2007, contains 10 synthetically generated classes summing up to 14 00014\,000 fault-free background textures and 2100 defectives, one for each background texture. In contrast, the more recent MIAD contains 105 000105\,000 images of various outdoor industrial maintenance scenarios. MIAD leverages 3D graphics software [20] to render realistic 3D scenes with varying surface textures, backgrounds and viewpoints on both fault-free and defective objects.
@@ 3 Industrial Screen Printing Anomaly Detection Dataset: ISP-AD
    High permitted sample variations arising from the sample itself, preprocessing, and imperfect imaging conditions.
@@ 3.1 Industrial Screen Printing and Defect Classes
The examined product is manufactured using a technique called screen or silk-screen printing [12]. It is a low-cost and highly automatable manufacturing process applied in various industries e.g. automotive, textile and electronics [67]. In a nutshell, during the printing process, ink is deposited through a stencil with a predefined design onto the front and/or backside of a polymer carrier foil. The subsequent repetition using different stencils and colors results in a multilayer decorated foil plate with a high-quality appealing design. However, the complex manufacturing process is error-prone at basically every step of production. This results in a wide range of different defect appearances and classes. The investigated defect classes within this publication are visualized in Fig. 1 and Fig. 2. They can be divided in punctual defects (Fig. 1), such as inclusions, scratches, dots, pinholes, printing or screen defects or area defects (Fig. 2), such as pattern misalignment, squeegee strokes and grid defects. As described in [39], three different optical modalities, consisting of darkfield illumination in Line Scan Modality 1 (LSM-1), brightfield illumination in Area Scan Modality (ASM), and transmission illumination in Line Scan Modality 2 (LSM-2), were mandatory to visualize the wide range of different defect classes.
For example, most defects in the transparent top layer, such as scratches or mechanical deformations, are only visible under brightfield illumination used in ASM, whereas printing or screen defects are captured under darkfield conditions by means of LSM-1. In LSM-2, transmission illumination is used to detect defects such as pinholes or inhomogeneities within the print layers. For a detailed description of the imaging modalities and a visualization of different defect classes and their root causes, we refer to sections 2 and 3 in [38].
In general, defects can be characterized as small (few px in extension in relation to large sized field of views (FOV) of up to ${10}^{6}\text{\,}{\mathrm{mm}}^{2}$, with minimum object pixel sizes of approximately 75 µm) and weakly contrasted. They are embedded within structured background patterns that exhibit high permissible variability, making visual separation difficult due to low contrast, shown by the fiber in Fig. 1. As visualized in Fig. 1 and Fig. 2, defect characteristics are strongly dependent on both the imaging modality and the specific defect class.
@@ 3.2 Dataset Generation
The generation of appropriate datasets builds the basis for achieving the required defect detection performance of machine learning based methods applied on the production line. Therefore, the design of validation and test splits mimicking the inspection process with its imbalanced data distributions is of utmost importance to perform reliable evaluations. Furthermore, generated data distributions must resemble the imperfect imaging conditions caused by environmental influences (e.g., varying illumination), data processing (varying background, viewpoint and segmentation artifacts), as well as permitted sample variations.
The proposed dataset is based on the work presented in [38], including its efficient data preprocessing and labeling workflow. All samples were collected directly from a real-world screen printing production line at a single industrial site, representing one product design. The investigated product is used in the automotive industry as a decorative element.
Prior to image acquisition, fault-free and defective samples were pre-sorted by domain experts. The sample state is characterized as fixed, thus additional temporal sample alterations can be excluded. Moreover, the samples were selected from different production batches, resembling the permitted design variability. Images were acquired by means of all three optical modalities (LSM-1, LSM-2, ASM) of the inspection system demonstrator installed on the factory floor. The acquired images of fault-free and defective samples were assigned to separate training and test splits. This step avoids any unwanted correlation of fault-free and defective patches during the subsequent patch extraction process. A detailed description of the utilized inspection system demonstrator, including its data processing steps, is provided in Section 4 of [38].
The extraction of augmented fault-free training patches (256 x 256 px) at random positions within the region of interest (ROI) is integrated into an automated procedure (steps 1 to 5 in Section 5.1 [38]).
Furthermore, following the central cropping in step 4, synthetic defects according to Section 5.2 are generated in 50 % of the fault-free patches. The utilized algorithm, proposed by [27], consists of four steps:
  1. 1.
    Generation of a binary defect skeleton, based on a stochastic process resembling a random walk with momentum.
  2. 2.
    Generation of a random defect texture.
  3. 3.
    Modification of the fault-free patch using the generated texture.
  4. 4.
    Final assessment of defect visibility and rejection of synthesized defects below the visibility threshold.
As described in Section 5.2 in [38], hyperparameters of the random variables were manually adjusted to synthesize bright- and dark-contrasted, punctuate and filamentous morphologies, aiming to imitate real-world defect characteristics, in all three optical modalities (Fig. 3). As a result, balanced supervised training datasets – also referred to as balanced synthetic streams – were obtained.
Sufficient defect visibility (e.g., via the sum of squared residuals of patch differences as proposed in [27]) was determined by manually inspecting up to 100 synthesized defects per data split by a domain expert, and adjusting defect-specific thresholds and hyperparameters accordingly, before applying automatic synthetization to several tens of thousands of patches. Due to compliance regulations, a detailed listing of hyperparameters and their selected values was omitted. However, a full description of the algorithm is available in [27].
@@
Above described workflow minimizes the manual labeling effort to known defective samples and regions. The underlying image-level labels can be considered "weak" since no ground truth masks are provided, thereby eliminating the need for elaborate pixel-level annotation. In general, patch-based processing is an efficient augmentation technique in upscaling available data, enabling extraction of several 100 000100\,000 patches from a few acquired samples with large FOVs. Theoretically, this automatic patch extraction process is unlimited, although it is restricted by available fault-free samples and FOV sizes [38].
@@ 3.2.1 Data Splits
The aim of this publication is to advance research in solving real-world industrial inspection problems by leveraging both synthetic and available real defects accumulated during production processes. Therefore, supervised and unsupervised datasets for each optical modality, acquired from the factory floor, have been created. The supervised training splits include up to approximately 100 000100\,000 augmented fault-free patches and 100 000100\,000 synthetic defective patches, forming a large-scale balanced synthetic training stream. Additionally, real defective patches, extracted from defective samples representing various defect classes, are categorized into two groups: "area" and "points". For unsupervised approaches, the fault-free training data includes up to 3678 patches in LSM-1, with reduced splits containing a maximum of 500 patches available for all modalities. These smaller splits are designed to minimize computational effort for e.g. memory demanding methods. The test data for both supervised and unsupervised approaches is identical, mimicking typical imbalanced inspection data distributions. Table 7 summarizes the proposed dataset with its synthetic and real defects available for supervision.
@@ 3.2.4 Additional Data
In addition to proposed data splits, supplementary fault-free patches for all modalities are provided. These patches are the unaugmented counterparts to the fault-free supervised training data described above. They consist of 20 00020\,000 and 10 00010\,000 fault-free patches in rgb format for LSM-1 and LSM-2, and 20 00020\,000 patches in single-channel format for ASM, all stored as .png files.
@@ 3.2.5 Limitations
General limitations include the presentation of a single product design due to compliance regulations, as well as a constrained set of defect classes. However, the observed defect diversity arises from process-related variability and is not strictly limited by the design itself. In addition, ground truth masks for real defective training data are not included, as the focus was set on weakly labeled training data conditions. Altough, their inclusion may be considered in future dataset revisions.
@@ 4 Defect Detecion Methods
The recently published work [38] introduced an efficient data preprocessing workflow utilizing weakly labeled defective data in a supervised training approach. Furthermore, the possible extension to a mixed training strategy using both synthetic and available real defective data [28] was outlined, enhancing defect detection sensitivity of overlooked defect classes.
Therefore, the evaluation of the mixed training strategy on the presented demanding imaging modalities should give further insights in leveraging supervised methods in industrial anomaly detection. For comparison several SOTA unsupervised anomaly detection methods are investigated. Utilized approaches as well as their training settings will be described in the following section.
@@ 4.1 Mixed Supervised Training
The proposed mixed supervised training strategy showed increased detection performance by leveraging both synthetic and real defects during training, compared to training separately on synthetic or real ones [28]. Thus, depending on the availability of real defects, the method can be applied in cold-start scenarios, starting with purely synthetic training and transitioning to a mixed setup as real defects accumulate. This is an important capability in industrial applications, as the model can be improved step by step with newly available defective samples.
However, the ability to synthesize certain defect classes using the algorithm in [27], such as squeegee strokes shown in Fig. 2, is limited or even infeasible. Nonetheless, such model-free approaches, as described in [77], do not rely on real defective samples for subsequent model training or conditioning, unlike generative models such as GAN- or diffusion-based methods [24, 32, 26]. They may therefore represent a more deployment-friendly strategy, particularly in cold-start scenarios where no real defective samples are available. In other words, model-free anomaly synthetization refers to methods that generate artificial defects through handcrafted, augmentation-based, or procedural rules without training a generative model [44, 84, 68, 27].
@@
Thus, even small amounts of previously unseen real defects may contribute to learning feature representations that could not be captured by synthetization or were absent in the available real data. Additionally, incorporating real defects, similar to synthesized ones, increases the feature diversity of already seen defects. Throughout this work, we use the term few-shot to refer specifically to training or defect modeling regimes with only a limited amount of real defective samples available (e.g., 10 patches), while the majority of training data is still provided by synthetic defective or fault-free patches (see Section 3.2.1 on data splits).
@@
Utilizing the above-introduced datasets, a DCNN pretrained on ImageNet [35] was trained in a supervised manner. Synthetic defective patches were generated in 50% of the fault-free patches.
During training, each patch in this balanced synthetic stream was replaced by an augmented real defective patch with an injection probability $p_{\text{inj}}$, chosen within the range
$p_{\text{inj}}\in\left[\tfrac{1}{B},\,0.25\right],$ (1)
and kept constant during each training stage, where $B$ denotes the batch size.
The injection probability $p_{\text{inj}}$ can be tuned to the quantity and diversity of available real defects: lower values favor specificity, while higher values improve sensitivity but also increase the risk of bias toward injected feature distributions. Therefore, the lower bound ensures that, in expectation, at least one real defect is injected into each batch, while the upper bound of 25% was chosen empirically to prevent real defects from dominating the synthetic stream.
Unlike dataset-level mixing approaches [58, 21] or deterministic batch-level schemes that enforce one real defect per batch [33], our method introduces a tunable stochastic injection. This allows adaptation to defect availability and synthesis capabilities while maintaining variability across training batches, ensuring that real defects complement rather than dominate the synthetic stream.
The proposed batch-wise stochastic injection scheme treats each patch independently. Consequently, the expected fraction of real defects in any batch equals $p_{\text{inj}}$, regardless of whether the underlying stream consists of purely fault-free patches or of a balanced synthetic–fault-free mixture.
This formulation highlights the connection to our previously published oversampling approach [38]: when the balanced synthetic stream is replaced by a pure fault-free stream and the injection probability is set to $p_{\text{inj}}=0.5$, the two approaches converge. In this special case, each batch contains, in expectation, 50% fault-free and 50% real defective patches.
Supervised binary classification is inherently a closed-set problem, in contrast to the open-set problem of anomaly detection described above. In production environments, such as the investigated screen printing process, possible defect causes are often known in advance, even if corresponding defective samples are not yet available. This effectively reduces the defect space to a restricted open-set problem, where not all defects are yet observed, but their characteristics are likely to fall within a range of expected or process-related visual features. To address this setting, our mixed supervised training can be characterized as an adaptable closed-set strategy, operating within this bounded defect space defined by process-specific visual characteristics, both observed and yet unobserved. Adapting the classifier to emerging, previously unseen defect classes with a minimal set of recently available samples would be an important capability in industrial settings.
@@
The following experiments in Section 5.2 aim to investigate the generalization capabilities of the proposed mixed training strategy for previously unseen defects (e.g., defects that could not be synthesized). In addition, its scalability is examined under varying proportions of collected real defects across all imaging modalities: LSM-1, LSM-2, and ASM. Before the main evaluations, ablation experiments across various backbones were performed.